# Wasserstein Transfer Learning

### v1

@@ Abstract
Transfer learning is a powerful paradigm for leveraging knowledge from source domains to enhance learning in a target domain. However, traditional transfer learning approaches often focus on scalar or multivariate data within Euclidean spaces, limiting their applicability to complex data structures such as probability distributions. To address this, we introduce a novel framework for transfer learning in regression models, where outputs are probability distributions residing in the Wasserstein space. When the informative subset of transferable source domains is known, we propose an estimator with provable asymptotic convergence rates, quantifying the impact of domain similarity on transfer efficiency. For cases where the informative subset is unknown, we develop a data-driven transfer learning procedure designed to mitigate negative transfer. The proposed methods are supported by rigorous theoretical analysis and are validated through extensive simulations and real-world applications.
@@ 1.2 Related Work
Transfer learning. Transfer learning aims to improve performance in a target population by leveraging information from a related source population and has seen wide application across domains (e.g., Huang et al. 2013, Gong et al. 2012, Choi et al. 2017, Zhou et al. 2024). Recent theoretical developments have focused on regression in Euclidean settings, including high-dimensional linear (Li et al. 2022) and generalized linear models (Tian & Feng 2023), nonparametric regression (Cai & Pu 2024, Lin & Reimherr 2024), and function mean estimation from discretely sampled data (Cai et al. 2024). In parallel, optimal transport has been used to measure distributional shifts for domain adaptation (Courty et al. 2017, Redko et al. 2017). However, to the best of our knowledge, no existing work has investigated transfer learning in regression models where outputs are probability distributions residing in the Wasserstein space. This represents a significant gap in the literature, highlighting the need for novel methodologies that address this challenging yet important setting.
@@ 2.1 Notations
[…]
@@ 2.2 Wasserstein Space
[…]
@@ 3.1 Setup
We consider a transfer learning problem where target data $\{(X_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}$ are sampled independently from the target population $(X^{(0)},\nu^{(0)})\sim\mathcal{F}_{0}$, and source data $\{(X_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}$ are sampled independently from the source population $(X^{(k)},\nu^{(k)})\sim\mathcal{F}_{k}$, for $k=1,\ldots,K$. The goal is to estimate the target model using both the target data and the source data from $K$ related studies.
For $k=0,\ldots,K$, assume $X^{(k)}$ has mean $\theta_{k}$ and covariance $\Sigma_{k}$, with $\Sigma_{k}$ positive definite. Define the empirical mean and covariance of $\{X_{i}^{(k)}\}_{i=1}^{n_{k}}$ as $\overline{X}_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}X_{i}^{(k)}$ and $\widehat{\Sigma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}(X_{i}^{(k)}-\overline{X}_{k})(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}.$ For a fixed $x\in\mathbb{R}^{p}$, the weight function is $s_{G}^{(k)}(x)=1+(X^{(k)}-\theta_{k})^{\mathrm{\scriptscriptstyle T}}\Sigma_{k}^{-1}(x-\theta_{k}),$ with the sample version $s_{iG}^{(k)}(x)=1+(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}\widehat{\Sigma}_{k}^{-1}(x-\overline{X}_{k}).$ The target regression function for a given $x\in\mathbb{R}^{p}$ is then $m_{G}^{(0)}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{G}^{(0)}(x)d_{\mathcal{W}}^{2}(\mu,\nu^{(0)})\}.$ In the following, we present details on transfer learning for global Fréchet regression, where the key difference in the local Fréchet regression setting is the use of a different weight function. The technical details for transfer learning in local Fréchet regression are therefore deferred to Appendix D.
@@ 3.2 Wasserstein Transfer Learning
We propose the Wasserstein Transfer Learning (WaTL) algorithm, which combines information from source datasets under the assumption that all source data are “informative enough.” This assumption implies that the discrepancies between the source and target are small enough to enhance estimation compared to using only the target. When this condition is met for all source datasets, the informative set is given by $\mathcal{A}_{\psi}=\{1,\ldots,K\}$. The detailed steps of WaTL are presented in Algorithm 1.
@@
In Step 1, the initial estimate $\widehat{f}$ aggregates information from both the target and source, weighted by their respective sample numbers.
While this step incorporates valuable auxiliary information, the resulting estimate may be biased due to distributional differences between the target and source populations. In Step 2, the bias in $\widehat{f}$ is corrected by focusing on the target data. The regularization term $\lambda\|g-\widehat{f}(x)\|_{2}$ ensures a balance between target-specific precision and auxiliary-informed robustness. Theoretical guidelines for selecting $\lambda$ are provided in Theorem 2. The final step projects the corrected estimate $\widehat{f}_{0}$ onto the Wasserstein space, ensuring the output $\widehat{m}^{(0)}_{G}(x)$ satisfies the intrinsic geometry of $\mathcal{W}$. Such a projection exists and is unique as $\mathcal{W}$ is a closed and convex subset of $L^{2}(0,1)$.
@@ 3.3 Adaptive Selection of Informative Sources
0:  Target and source data $\{(x_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}\cup\big{(}\cup_{1\leq k\leq K}\{(x_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}\big{)}$, regularization parameter $\lambda$, number of informative sources $L$, and query point $x\in\mathbb{R}^{p}$.
@@
1:  Compute discrepancy scores. For each source dataset $k=1,\ldots,K$, compute the empirical discrepancy: $\widehat{\psi}_{k}=\|\widehat{f}^{(0)}(x)-\widehat{f}^{(k)}(x)\|_{2},$ where $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F^{-1}_{\nu_{i}^{(k)}}$. Construct the adaptive informative set by selecting the $L$ smallest discrepancy scores $\widehat{\mathcal{A}}=\{1\leq k\leq K:\widehat{\psi}_{k}\text{ is among the smallest }L\text{ values}\}.$
2:  Weighted auxiliary estimator: $\widehat{f}(x)=\frac{1}{\sum_{k\in\widehat{\mathcal{A}}\cup\{0\}}n_{k}}\sum_{k\in\widehat{\mathcal{A}}\cup\{0\}}n_{k}\widehat{f}^{(k)}(x).$
@@ 4 Theory
For $k=0,\ldots,K$, the covariate $X^{(k)}$ is subguassian with $\|X^{(k)}\|_{\Psi_{2}}\in[\sigma_{1},\sigma_{2}]$, the mean vector satisfies $\|\theta_{k}\|\leq R_{1}$, and the spectrum of the covariance matrix $\Sigma_{k}$ lies within the interval $[R_{2},R_{3}]$. Moreover, $\nu^{(k)}$ is supported on a bounded interval.
@@
Let $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F_{\nu_{i}^{(k)}}^{-1}$ and its population counterpart $f^{(k)}(x)=E\{s_{G}^{(k)}(x)F_{\nu^{(k)}}^{-1}\}$ for $k=0,\ldots,K$. Then under Condition 1, $\|\widehat{f}^{(k)}(x)-f^{(k)}(x)\|_{2}=O_{p}(n_{k}^{-1/2})$.
@@
Theorem 2 can be contrasted with the convergence rate of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, where the rate for regression with distributional outputs is $d_{\mathcal{W}}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}(n_{0}^{-1/2})$. The WaTL algorithm achieves a faster convergence rate when there are sufficient source data and the auxiliary sources are informative enough, satisfying $\psi\lesssim n_{0}^{-1/2-\epsilon}$. This result highlights that knowledge transfer from auxiliary samples can significantly enhance the learning performance of the target model, provided the auxiliary sources are closely aligned with the target data.
@@
Condition 3 ensures that the source datasets can be effectively partitioned into two groups: one that is informative and another that is sufficiently different from the target data. This separation guarantees that the informative set can be accurately identified. Under this condition, setting the number of informative sources as $L=|\mathcal{A}|$, we establish the following rate of convergence for the AWaTL algorithm. In practice, $L$ can be decided by cross-validation.
@@ 5 Numerical Experiments
In this section, we evaluate the performance of the proposed WaTL algorithm, alongside two baseline approaches: the global Fréchet regression using only target data (Only Target) and using only source data (Only Source). Consider $K=5$ source sites. The data is generated as follows. For the target population, we sample $X^{(0)}\sim{\rm U}(0,1)$ and generate the response distribution, represented by its quantile function, as
@@
where $Z^{(0)}\sim N(0.5,1)|_{(0,1)}$ and $w^{(0)}\sim N(0,1)|_{(-0.5,0.5)}$. Here, $N(\mu,\sigma^{2})|_{(a,b)}$ denotes a normal distribution with mean $\mu$ and variance $\sigma^{2}$, truncated to the interval $(a,b)$. For source populations, we define $\psi_{k}=0.1k$ for $k=1,\ldots,K$ and generate $X^{(k)}\sim{\rm U}(0,1)$. The corresponding response distribution is generated as
@@
where $Z^{(k)}\sim N(0.5,1-\psi_{k})|_{(0,1)}$ and $w^{(k)}\sim N(0,1)|_{(-0.5,0.5)}$. Consequently, for each predictor $x$, the true regression function is $m_{G}^{(k)}(x)=(1-x)u+xF^{-1}_{Z^{(k)}}(u)$ for $k=0,1,\ldots,K$.
We vary the target sample size $n_{0}$ from $200$ to $800$, while the source sample size is set as $n_{k}=k\tau$, where $\tau\in\{100,200\}$ and $k=1,\ldots,K$. The regularization parameter $\lambda$ in Algorithm 1 is selected via five-fold cross-validation, varying from 00 to $3$ in increments of $0.1$. To evaluate performance, we sample $100$ predictors uniformly from the target distribution. Using Algorithm 1, we compute $\widehat{m}_{G}^{(0)}(x)$ and compare it with the corresponding estimates obtained from global Fréchet regression using only target or source data. Performance is assessed using the root mean squared prediction risk $\mathrm{RMSPR}=\sqrt{\frac{1}{100}\sum_{i=1}^{100}d_{\mathcal{W}}^{2}\big{(}\widehat{m}_{G}^{(0)}(x_{i}),{m}_{G}^{(0)}(x_{i})\big{)}}$, where $x_{i}$ denotes the sampled predictor, $\widehat{m}_{G}^{(0)}(x_{i})$ is the estimated function, and ${m}_{G}^{(0)}(x_{i})$ represents the ground truth. To ensure robustness, we repeat the simulation $50$ times and report the average RMSPR, as shown in Figure 1 (a).
@@
Comparing the two panels of Figure 1 (a), we also observe that WaTL improves as the source sample size increases, confirming its ability to effectively integrate source information. This demonstrates the benefit of multi-source transfer learning, where WaTL balances knowledge from both target and source domains to achieve improved prediction.
@@
We further evaluate the effectiveness of AWaTL in selecting informative sources. In this experiment, we set $L=2$ and $n_{k}=100$ for $k=0,\ldots,5$. The similarity parameters are set as $\psi_{k}=0.1$ for $k=1,2$ (informative sources), and $\psi_{k}$ from $0.2$ to $1$ in increments of $0.1$ for $k=3,4,5$ (uninformative sources). We repeat each setting 100 times and report the selection rates in Figure 1 (b). The results show that AWaTL successfully identifies informative sources, with selection rates for sources $1$ and $2$ rapidly increasing and reaching perfect accuracy when $\psi>0.6$. This highlights the robustness of AWaTL in distinguishing useful sources under varying similarity levels.

### v2

@@ Abstract
Transfer learning is a powerful paradigm for leveraging knowledge from source domains to enhance learning in a target domain. However, traditional transfer learning approaches often focus on scalar or multivariate data within Euclidean spaces, limiting their applicability to complex data structures such as probability distributions. To address this limitation, we introduce a novel transfer learning framework for regression models whose outputs are probability distributions residing in the Wasserstein space. When the informative subset of transferable source domains is known, we propose an estimator with provable asymptotic convergence rates, quantifying the impact of domain similarity on transfer efficiency. For cases where the informative subset is unknown, we develop a data-driven transfer learning procedure designed to mitigate negative transfer. The proposed methods are supported by rigorous theoretical analysis and are validated through extensive simulations and real-world applications. The code is available at https://github.com/h7nian/WaTL.
@@ 1.2 Related Work
Transfer learning. Transfer learning aims to improve performance in a target population by leveraging information from a related source population and has seen wide application across domains (e.g., Huang et al. 2013, Gong et al. 2012, Choi et al. 2017, Zhou, Liu, Li & Cai 2024, Zhou, Li, Cai & Liu 2024). Recent theoretical developments have focused on regression in Euclidean settings, including high-dimensional linear (Li et al. 2022) and generalized linear models (Tian & Feng 2023), nonparametric regression (Cai & Pu 2024, Lin & Reimherr 2024), and function mean estimation from discretely sampled data (Cai et al. 2024). In parallel, optimal transport has been used to measure distributional shifts for domain adaptation (Courty et al. 2017, Redko et al. 2017). However, to the best of our knowledge, no existing work has investigated transfer learning in regression models where outputs are probability distributions residing in the Wasserstein space. This represents a significant gap in the literature, highlighting the need for novel methodologies that address this challenging yet important setting.
@@ 2.1 Notations
Let $L^{2}(0,1)$ be the space of square-integrable functions over the interval $(0,1)$, with the associated $L^{2}$ norm and metric denoted by $\|\cdot\|_{2}$ and $d_{L^{2}}$, respectively. To be specific, $\|g\|_{2}=(\int_{0}^{1}g^{2}(z)dz)^{1/2}\text{ and }d_{L^{2}}(g_{1},g_{2})=\|g_{1}-g_{2}\|_{2}$. For a vector $Z$, $\|Z\|$ denotes the Euclidean norm. Given a matrix $\Sigma$, we define its spectrum as the set of its singular values. For a sub-Gaussian random vector $X$, we define the sub-Gaussian norm as $\|X\|_{\Psi_{2}}:=\sup_{\|v\|=1}\inf\left\{t>0:E(e^{\langle X,v\rangle^{2}/t^{2}})\leq 2\right\}.$
@@ 2.2 Wasserstein Space
Let $\mathcal{W}$ denote the space of probability distributions on $\mathbb{R}$ with finite second moments, equipped with the 2-Wasserstein, or simply Wasserstein, metric. For two distributions $\mu_{1},\mu_{2}\in\mathcal{W}$, the Wasserstein metric is given by $d_{\mathcal{W}}^{2}(\mu_{1},\mu_{2})=\inf_{\pi\in\Pi(\mu_{1},\mu_{2})}\int_{\mathbb{R}\times\mathbb{R}}|s-t|^{2}\,d\pi(s,t)$, where $\Pi(\mu_{1},\mu_{2})$ denotes the set of all joint distributions with marginals $\mu_{1}$ and $\mu_{2}$ (Kantorovich 1942). For a probability measure $\mu\in\mathcal{W}$ with cumulative distribution function $F_{\mu}$, we define the quantile function $F_{\mu}^{-1}$ as the left-continuous inverse of $F_{\mu}$, such that $F_{\mu}^{-1}(u)=\inf\{t\in\mathbb{R}|F_{\mu}(t)\geq u\}$, for $u\in(0,1)$. It has been established (Villani 2009) that the Wasserstein metric can be expressed as the $L^{2}$ metric between quantile functions:
$d_{\mathcal{W}}^{2}(\mu_{1},\mu_{2})=\int_{0}^{1}\left\{F_{\mu_{1}}^{-1}(u)-F_{\mu_{2}}^{-1}(u)\right\}^{2}du.$ (1)
@@ 3.1 Setup
We consider a transfer learning problem where target data $\{(X_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}$ are sampled independently from the target population $(X^{(0)},\nu^{(0)})\sim\mathcal{F}_{0}$, and source data $\{(X_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}$ are sampled independently from the source population $(X^{(k)},\nu^{(k)})\sim\mathcal{F}_{k}$, for $k=1,\ldots,K$. The goal is to estimate the target model using both the target data and source data from $K$ related studies.
For $k=0,\ldots,K$, assume $X^{(k)}$ has mean $\theta_{k}$ and covariance $\Sigma_{k}$, with $\Sigma_{k}$ positive definite. Define the empirical mean and covariance of $\{X_{i}^{(k)}\}_{i=1}^{n_{k}}$ as $\overline{X}_{k}=n_{k}^{-1}\sum_{i=1}^{n_{k}}X_{i}^{(k)}$ and $\widehat{\Sigma}_{k}=\frac{1}{n_{k}}\sum_{i=1}^{n_{k}}(X_{i}^{(k)}-\overline{X}_{k})(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}.$ For a fixed $x\in\mathbb{R}^{p}$, the weight function is $s_{G}^{(k)}(x)=1+(X^{(k)}-\theta_{k})^{\mathrm{\scriptscriptstyle T}}\Sigma_{k}^{-1}(x-\theta_{k}),$ with the sample version $s_{iG}^{(k)}(x)=1+(X_{i}^{(k)}-\overline{X}_{k})^{\mathrm{\scriptscriptstyle T}}\widehat{\Sigma}_{k}^{-1}(x-\overline{X}_{k}).$ The target regression function for a given $x\in\mathbb{R}^{p}$ is then $m_{G}^{(0)}(x)=\operatorname*{arg\,min}_{\mu\in\mathcal{W}}E\{s_{G}^{(0)}(x)d_{\mathcal{W}}^{2}(\nu^{(0)},\mu)\}.$ In the following, we present details on transfer learning for global Fréchet regression, where the key difference in the local Fréchet regression setting is the use of a different weight function. The technical details for transfer learning in local Fréchet regression are therefore deferred to Appendix 11.
@@ 3.2 Wasserstein Transfer Learning
We propose the Wasserstein Transfer Learning (WaTL) algorithm, which combines information from source datasets under the assumption that all source data are informative enough. This assumption implies that the discrepancies between the source and target are small enough to enhance estimation compared to using only the target. When this condition is met for all source datasets, the informative set is given by $\mathcal{A}_{\psi}=\{1,\ldots,K\}$. The detailed steps of WaTL are presented in Algorithm 1.
@@
In Step 1, the initial estimate $\widehat{f}$ aggregates information from both the target and source, weighted by their respective sample sizes. While this step incorporates valuable auxiliary information, the resulting estimate may be biased due to distributional differences between the target and source populations. In Step 2, the bias in $\widehat{f}$ is corrected by focusing on the target data. The regularization term $\lambda\|g-\widehat{f}(x)\|_{2}$ ensures a balance between target-specific precision and auxiliary-informed robustness. Theoretical guidelines for selecting $\lambda$ are provided in Theorem 2. The final step projects the corrected estimate $\widehat{f}_{0}$ onto the Wasserstein space, ensuring the output $\widehat{m}^{(0)}_{G}(x)$ respects the intrinsic geometry of $\mathcal{W}$. This projection exists and is unique because $\mathcal{W}$ is a closed and convex subset of $L^{2}(0,1)$.
@@ 3.3 Adaptive Selection of Informative Sources
0: Target and source data $\{(x_{i}^{(0)},\nu_{i}^{(0)})\}_{i=1}^{n_{0}}\cup\big(\cup_{1\leq k\leq K}\{(x_{i}^{(k)},\nu_{i}^{(k)})\}_{i=1}^{n_{k}}\big)$, regularization parameter $\lambda$, prespecified number of informative sources $L$, and query point $x\in\mathbb{R}^{p}$.
@@
[…]
@@ 4 Theory
For $k=0,\ldots,K$, the covariate $X^{(k)}$ is sub-Gaussian with $\|X^{(k)}\|_{\Psi_{2}}\in[\sigma_{1},\sigma_{2}]$, the mean vector satisfies $\|\theta_{k}\|\leq R_{1}$, and the spectrum of the covariance matrix $\Sigma_{k}$ lies within the interval $R_{2},R_{3}]$. Moreover, $\nu^{(k)}$ is supported on a bounded interval.
@@
Let $\widehat{f}^{(k)}(x)=n_{k}^{-1}\sum_{i=1}^{n_{k}}s_{iG}^{(k)}(x)F_{\nu_{i}^{(k)}}^{-1}$ and its population counterpart be defined as $f^{(k)}(x)=E\{s_{G}^{(k)}(x)F_{\nu^{(k)}}^{-1}\}$ for $k=0,\ldots,K$. Then under Condition 1, $\|\widehat{f}^{(k)}(x)-f^{(k)}(x)\|_{2}=O_{p}(n_{k}^{-1/2})$.
@@
Theorem 2 can be compared to the convergence rate of global Fréchet regression (Petersen & Müller 2019) applied solely to the target data, for which the rate is $d_{\mathcal{W}}(\widehat{m}_{G}^{(0)}(x),m_{G}^{(0)}(x))=O_{p}(n_{0}^{-1/2})$. The WaTL algorithm achieves a faster convergence rate when there are sufficient source data and the auxiliary sources are informative enough, satisfying $\psi\lesssim n_{0}^{-1/2-\epsilon}$. This result highlights that knowledge transfer from auxiliary samples can significantly enhance the learning performance of the target model, provided the auxiliary sources are closely aligned with the target data.
@@
Condition 3 ensures that the source datasets can be effectively partitioned into two groups: informative ones and those sufficiently different from the target data. This separation guarantees that the informative set can be accurately identified. Under this condition, by setting the number of informative sources to $L=|\mathcal{A}|$, we establish the following rate of convergence for the AWaTL algorithm. In practice, $L$ can be decided by cross-validation.
@@
While the above analysis assumes that each probability distribution is fully observed, an assumption commonly adopted in the distributional data literature (Petersen & Müller 2019, Chen et al. 2023), real-world applications often provide only independent samples drawn from underlying distributions. In such cases, this limitation can be overcome by replacing unobservable distributions with their empirical counterparts, constructed from sample observations (Zhou & Müller 2024). Additional details and theoretical justification for this extension are provided in Appendix 12.
@@ 5 Numerical Experiments
In this section, we evaluate the performance of the proposed WaTL algorithm, alongside two baseline approaches: the global Fréchet regression using only target data (Only Target) and using only source data (Only Source). Consider $K=5$ source sites. The data are generated as follows. For the target population, we sample $X^{(0)}\sim{\rm U}(0,1)$ and generate the response distribution, represented by its quantile function, as
@@
where $Z^{(0)}\sim N(0.5,1)|_{(0,1)}$ and $w^{(0)}\sim N(0,1)|_{(-0.5,0.5)}$. Here, $N(\mu,\sigma^{2})|_{(a,b)}$ denotes a normal distribution with mean $\mu$ and variance $\sigma^{2}$, truncated to the interval $(a,b)$. For source populations, we define $\psi_{k}=0.1k$ for $k=1,\ldots,K$, and generate $X^{(k)}\sim{\rm U}(0,1)$. The corresponding response distribution is generated as
@@
where $Z^{(k)}\sim N(0.5,1-\psi_{k})|_{(0,1)}$ and $w^{(k)}\sim N(0,1)|_{(-0.5,0.5)}$. Consequently, for each predictor $x$, the true regression function is $m_{G}^{(k)}(x)=(1-x)u+xF^{-1}_{Z^{(k)}}(u)$, for $k=0,1,\ldots,K$.
We vary the target sample size $n_{0}$ from $200$ to $800$, while the source sample size is set as $n_{k}=k\tau$, where $\tau\in\{100,200\}$ and $k=1,\ldots,K$. The regularization parameter $\lambda$ in Algorithm 1 is selected via five-fold cross-validation, ranging from 0 to $3$ in increments of $0.1$. To evaluate performance, we sample $100$ predictors uniformly from the target distribution. Using Algorithm 1, we compute $\widehat{m}_{G}^{(0)}(x)$ and compare it with the corresponding estimates obtained from global Fréchet regression using only target or source data. Performance is assessed using the root mean squared prediction risk $\mathrm{RMSPR}=\sqrt{\frac{1}{100}\sum_{i=1}^{100}d_{\mathcal{W}}^{2}\big(\widehat{m}_{G}^{(0)}(x_{i}),{m}_{G}^{(0)}(x_{i})\big)}$, where $x_{i}$ denotes the sampled predictor, $\widehat{m}_{G}^{(0)}(x_{i})$ is the estimated function, and ${m}_{G}^{(0)}(x_{i})$ represents the ground truth. To ensure robustness, we repeat the simulation $50$ times and report the average RMSPR, as shown in Figure 1(a).
@@
Comparing the two panels of Figure 1(a), we also observe that WaTL improves as the source sample size increases, confirming its ability to effectively integrate information from source domains. This demonstrates the benefit of multi-source transfer learning, where WaTL balances knowledge from both target and source domains to achieve improved prediction.
@@
We further evaluate the effectiveness of AWaTL in selecting informative sources. In this experiment, we set $L=2$ and $n_{k}=100$ for $k=0,\ldots,5$. The similarity parameters are specified as $\psi_{k}=0.1$ for $k=1,2$ (informative sources), and $\psi_{k}=\psi$, increasing from $0.2$ to $1$ in increments of $0.1$, for $k=3,4,5$ (uninformative sources). Each configuration is repeated $100$ times, and the corresponding selection rates are reported in Figure 1(b). The results show that AWaTL successfully identifies informative sources, with selection rates for sources $1$ and $2$ rapidly increasing and reaching perfect accuracy once $\psi>0.6$. These findings demonstrate the robustness of AWaTL in distinguishing useful sources under varying similarity levels.